OpenStudy (anonymous):

If 4.2 moles of copper metal reacts with 6.3 moles of silver nitrate, how many moles of silver metal can be formed, and how many moles of the excess reactant will be left over when the reaction is complete? Unbalanced equation: Cu + AgNO3 → Cu(NO3)2 + Ag Be sure to show all of your work. @taramgrant0543664 @photon336

OpenStudy (anonymous):

xCu + 2xAgNO3 --> xCu(NO3)2 + 2xAg If 2X=6.3, find X and hence no of moles reacted nd unreacted

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Here AgNO3 is the limiting reagent, so you have to calculate based on it

OpenStudy (anonymous):

Can you walk me through this whole equation because it's so confusing

OpenStudy (anonymous):

The balanced eqn says that 1 mole of Cu requires 2 moles of AgNO3 to react with it. Since, you have 4.2 moles of Cu metal, it means you require 8.4 moles of AgNO3, which you don't have. So, you have to consider that you have 6.3 moles of AgNO3 that reacts with 6.3/2 moles(=3.15 moles) of Cu. So, no of moles of Ag formed=6.3 moles and, no of moles of unreacted Cu=4.2-3.15=1.05 moles

OpenStudy (anonymous):

Whats the balanced equation?

OpenStudy (anonymous):

Cu + 2AgNO3 --> Cu(NO3)2 + 2Ag

OpenStudy (anonymous):

Ok then what?

OpenStudy (anonymous):

The balanced eqn says that 1 mole of Cu requires 2 moles of AgNO3 to react with it. Since, you have 4.2 moles of Cu metal, it means you require 8.4 moles of AgNO3, which you don't have. So, you have to consider that you have 6.3 moles of AgNO3 that reacts with 6.3/2 moles(=3.15 moles) of Cu. So, no of moles of Ag formed=6.3 moles and, no of moles of unreacted Cu=4.2-3.15=1.05 moles

OpenStudy (kash_thesmartguy):

This link will help you - http://openstudy.com/study#/updates/53c1c134e4b00f624a91b3df